The meeting commenced at 1040 hours with the following in attendance:
|Mick Day ( Chairman )||-|
|Boyd Potts ( Secretary )||Orpheus C.C.|
|Sue Mabbett (Access Officer)||South Wales C.C.|
|Rhian Hicks ( Biological Recorder)||-|
|Ben Lovett ( Conservation Officer)||Morgannwg C.C.|
|Huw Jones||Draenan Diggers.|
|Martin Laverty||Grwp Ogofeydd Garimpeiros.|
|John Cooper||Chelsea S.S.|
|Chris Densham ( Trustee )||Oxford University C.C.|
|Charles Bailey (Trustee )||-|
Confirmed the meeting was quorate and opened the proceedings.
Ali Garman ( Survey Officer), Elsie Little ( CCC ), Spencer Drew (Treasurer), Pete Bolt ( BEC ), Andy Farrant ( Geological Recorder), Tim Long (Trustee - Now residing in the USA).
Noted the attendance of M. Laverty at the meeting and agreed the proposal under item 8 was correctly recorded. It was agreed unanimously, 'That the minutes be accepted as a true record'.
Having been away for nearly four months there was nothing to report except items already listed on the agenda.
Not able to attend but had forwarded accounts which were circulated at the meeting.
Noted that we still expect some Legal Fees in the near future but these should only amount to around to around pounds 200. There was a discussion on the accounts. Agreed we must make sure that our income covers running costs and it was accepted that they do so at present.
Not yet in the accounts is pound 80 Key deposit money paid in very recently. Proposed and accepted that the pound 100 itemised in the accounts be paid to the GCRT. It was then proposed by JC and seconded by BL: 'That the Treasurer's Report be accepted'. All agreed.
A report was circulated at the meeting.
SM stated that the numbers listed are probably lower than they should be due to problems over storage of the book. Some entries are illegible and others missing due to not finding the book. BL stated that Brian at the 'Lamb and Fox' is willing to have the book installed outside the pub. Agreed this needs to be implemented as soon as possible.
The padlock code has been changed and it was agreed that we may need to replace padlocks annually due to deterioration of same. Request made by Hereford CC for a permanent key. Discussion on this and their request was finally accepted provided we receive a letter of intent and the pound 20 fee.
The gate opening mechanism has been reported as being difficult to move. SM and CD will remedy this with the appropriate lubrication. Thanks were recorded to SM for all her work.
i. BL presented a Draft 'Taping Guidelines' to the meeting. This has been prepared to explain the reasons behind the present taping policy. Proposed by SM and seconded by JC; 'That the Taping Guidelines be adopted'. All agreed.
Request from Tim Guilford that he be allowed to co-ordinate a bat survey; mainly on the surface but with some underground work. The meeting stated that it had no objections so long as it is in cooperation with other bat researchers and he keeps the Biological Recorder, the Landowner and the Meeting informed where necessary.
ii. The Conservation Officer's Report was noted and the replacement of the Slaughter Canyon tape was supported. Tape stockpiles are being used. CB will supply more if requested.
More discussion on taping. Suggested a brief notice be placed in the entrance series explaining the reasons for Conservation taping. Also put taping information on the web site.
Agreed that data be supplied to Steve Roberts via RH for the web site. RH will endeavour to circulate data, via E. Mail, to the committee for preview and then liaise with SR and act as author for the data.
iii. CB questioned how GCRT are to deal with possible 2nd entrance access. Noted that they have key supplied via I.W-J. Agreed it has to be left to the Rescue Controller to assess whether access needed via that entrance. Also noted that GCRT need to be kept informed of any change of status of the entrances. There was a long discussion on these points.
Not in attendance.
BL commented that all available Grade2 data has been passed to SR for the web site. There are some gaps in the recorded data particularly as some Grade2 surveys have not yet been passed to the Group for inclusion. It was suggested that downloadable files be put on the web. Comment made that all available Grade2 data has been passed to the Grade5 surveyors.
Discussion on how long we wait for Grade2 information to be supplied but no decision was made on this point.
Long discussion held on survey data in general and the possibility of a national survey archive.
Noted that the Grade5 surveyors have done an excellent job and should be praised by all for their efforts.
i. RH investigated and commented on our Bat Policy after a request from the previous meeting. Noted that our present Code of Conduct does give guidelines for Bat protection. Suggestion made that we add 'explosives' to section 7. There was a long discussion on this. Other cave guidelines were studied; it was accepted we do not want to legislate too tightly. Finally agreed that we leave things as they are but bear in mind that explosives are implicit within the guidelines.
ii. Concern was expressed that specimens from the cave are not being recorded and are being dispersed and becoming lost or inaccessible. RH needs to be informed of specimens so they can be recorded.
Not in attendance. It was requested that AF be asked if he could supply the meeting with any information about his Rolex grant application. The meeting expressed its interest and willingness to help. ** (AF later supplied a report to the secretary. See article attached to the end of the Minutes.)
No new applications.
It was noted that some member clubs have not had a representative at the meeting for over two years.
Licence is now organised but still awaiting return from signing by the third Trustee (TL).
** ( Now signed and back with solicitors.) Agreed we send a gift and letter of thanks to Ursula Collie and her colleagues. CD to action. Brief discussion on the detail of the Licence. Noted that we need an inventory of equipment in the cave for which the Group accepts responsibility.
Path needs some repair. There was a discussion on the best approach to this. Noted that it has partially stabilised but that parts of the route do need improving and will be looked at in detail to determine the best method of repair.
Stile over the fence to be left until new fence is installed.
8(a)There is a film crew in the area. They may request a trip into the cave and if necessary this will be dealt with by members in the locality.
8(b)The SSSI status. The process is ongoing but will occur sometime in the future!
8(c)We need to obtain information on the status of the area as a RIG site?
8(d)Suggested by ML that the Minutes be circulated by E-Mail. BP agreed to do this where possible.
This has been arranged for Sunday 18th June 2000. To be held at the Gwesty Bach, Brynmawr, at 10.30am.